OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, 6) and has a slope of –1? x + y = –4 x + y = 4 x –y = –4 x –y = 4

OpenStudy (anonymous):

need help cant seem to figure it out :(

OpenStudy (anonymous):

What grade is this?

OpenStudy (anonymous):

im in 9th grade freshman class

OpenStudy (anonymous):

Oh, I'm sorry I can't help I'm only in 7th grade and I'm dumb at math

OpenStudy (mathmale):

You are given a point through which your line passes, as well as the slope of the line. Use the "point-slope formula for the equation of a straight line:" \[y-y _{1}=m(x-x _{1}).\]

OpenStudy (mathmale):

The given point is (-2,6). Thus, let x1=-2 and y1=6. Also substitute m=-1.

OpenStudy (mathmale):

that's the main work. The "standard form of an equation of a str. line" is Ax+By+C=0 or Ax+By=D. Once you have your first equation, re-write it in the above form to finish this problem.

OpenStudy (anonymous):

Okay thank you! Ill write it down now

OpenStudy (anonymous):

so y-6=-1x-(-2) would that be were I would start and just continue solving from there?